0.3x-0.2y=4, 0.4x+0.5y= -55/17, What is the solution of the system. Use integers or fractions.
Answer by checkley77(12844)

0.3x-0.2y=4 multiply by 5
0.4x+0.5y=-55/17 multiply by 2 & add.
1.5x-y=20
.8x+y=110/17
--------------------
2.3x=20+110/17
2.3x=(20*17+110)/17
2.3x=(340+110)/17
2.3x=450/17 cross multiply.
2.3x*17=450
39.1x=450
x=450/39.1
x=11.51 ans
.3*11.51-.2y=4
3.453-.2y=4
-.2y=4-3.453
-.2y=.547
y=.547/-.2
y=-2.735 ans.
Proof:
.4*11.51+.5*-2.735=-55/17
4.604-1.367=-55/17
-3.236~-3.235